The velocity of a moving vehicle is given by the equation v=(2t+3)^4 .Use the Chain Rule to determine an equation for the acceleration when a=dv/dt.
1
Expert's answer
2020-05-07T18:52:41-0400

We will use the Chain Rule for the following functions: "v = (2t+3)^4 = f(g(t))", where "f(t) = t^4, g(t) = 2t + 3" .

Since "\frac{df}{dt} = 4t^3" and "\frac{dg}{dt} = 2" , by the chain rule, we obtain that:


"a = \frac{dv}{dt} = \frac{df}{dg} \frac{dg}{dt} = 4(2t+3)^3 \cdot 2 = 8(2t+3)^3"